Bulah invested $2000 altogether. She made a profit of 15% on one investment, but lost 3% on the other investment. She ended up with a gain of $84. how much of the $2000 was in each of these investments?
Answer by checkley75(3666)

.15X-.03(2000-X)=84
.15X-60+.03X=84
.18X=84+60
.18X=144
X=144/.18
X=800 AMOUNT INVESTED @ 15%
2000-800=1200 INVESTED & LOST 3%
PROOF
800*.15-1200*.03=84
120-36=84
84=84